et X and Y be continuous random variables having joint density functio


1
Expert's answer
2021-09-28T20:38:10-0400

To find the constant c, we write

"1=∫^∞_{−∞}∫^∞_{−∞}f_{XY}(x,y)dxdy\\
1=∫^1_0∫^{1−x}_0cx+1dydx\\
1=∫^1_0(cx+1)(1−x)dx\\
1=\frac{1}{2}+\frac{1}{6}c."

Thus, we conclude c=3